# Federated Learning for Smart Grid: A Survey on Applications and Potential Vulnerabilities

### v1

@@ 3.3. Consumption Stage
In the area of optimal power usage (Rajesh et al., 2024; Abdulla et al., 2024), FL facilitates the extraction of typical electricity consumption patterns from smart meter data, enabling more efficient grid management. For load forecasting (Badr et al., 2023; Liu et al., 2023; Fekri et al., 2023), FL-based methods have been developed to enhance the accuracy of short-term residential load predictions by utilizing distributed data from multiple sources. In the case of electricity theft detection (Li et al., 2024a; Alshehri et al., 2024; Zafar et al., 2023), frameworks like FedDetect employ federated learning to effectively identify fraudulent activities while preserving consumer privacy.
@@
Optimal Power Usage Rajesh et al. (Rajesh et al., 2024) propose a FL framework designed to enhance privacy while optimizing power usage in SGs. The framework allows for collaborative learning across multiple participants without compromising individual user data. By securing power traces and enabling personalized recommendations for power usage optimization, the approach significantly improves privacy preservation in SG data analytics. This personalized aspect ensures that recommendations are tailored to the specific needs of users while maintaining data confidentiality. Similarly, Abdulla et al. (Abdulla et al., 2024) present an adaptive FL approach for energy consumption forecasting using smart meter data. Their method improves forecasting accuracy by utilizing smart meter data without sharing raw information, ensuring user privacy. The adaptive nature of the approach allows it to adjust to individual consumption patterns, providing personalized forecasting models for each participant. This enhances both the accuracy of energy predictions and the privacy of users in the SG system.
@@
Load Forecasting Badr et al. (Badr et al., 2023) introduce a privacy-preserving and communication-efficient energy prediction scheme based on FL for SGs. Their approach enables accurate load forecasting while maintaining data confidentiality, which is crucial for infrastructure planning and power dispatching. By enhancing communication efficiency in the FL process, this method helps reduce power outages and equipment failures, making it an important tool for SG operations. Liu et al. (Liu et al., 2023) propose FedForecast, a FL framework designed for short-term probabilistic individual load forecasting. This framework provides comprehensive information about future load uncertainties while addressing data isolation and privacy concerns inherent in traditional centralized forecasting methods. By maintaining data confidentiality across households or smart meters, FedForecast improves the accuracy and reliability of short-term load forecasts. Fekri et al. (Fekri et al., 2023) develop an asynchronous adaptive FL algorithm for distributed load forecasting with smart meter data. This approach enhances the scalability and efficiency of load forecasting in SGs by adapting asynchronously to the data, thus addressing issues of privacy, security, and communication overhead found in traditional centralized methods. By doing so, the algorithm ensures accurate load forecasting while preserving the privacy of distributed data.
@@
Energy Theft Detection Li et al. (Li et al., 2024a) introduce FUSE, a framework that combines FL with U-shape split learning for detecting electricity theft. FUSE identifies abnormal electricity consumption behaviors while preserving the privacy of both user data and the detection model. By integrating FL with split learning, FUSE enhances data protection and model security in smart grid systems, offering a robust solution for electricity theft detection. Alshehri et al. (Alshehri et al., 2024) develop a deep anomaly detection framework that utilizes FL to detect electricity theft, including zero-day threats, in smart power grids. This framework preserves privacy while effectively identifying new and previously unseen attacks, thereby enhancing the security and stability of SGs. By addressing the challenge of detecting sophisticated electricity theft attempts, the framework plays a critical role in maintaining grid reliability. Zafar et al. (Zafar et al., 2023) present a FL-assisted hybrid deep learning model for electricity theft detection using smart meter data. This model enhances the security and reliability of SGs, particularly in the context of Industry 5.0. By leveraging FL to keep data decentralized, the model addresses privacy concerns while effectively detecting electricity theft. This hybrid approach offers a comprehensive solution to safeguard SGs from fraudulent activities.
@@ 4.1. Vulnerabilities in Generation Stage
In the context of SG, generation refers to the production of electricity from various sources, including both traditional and renewable energy (Tuballa and Abundo, 2016). It is a crucial component of the SG pipeline, serving as the starting point for the flow of electricity throughout the entire system. As generation systems become more integrated and digitalized, privacy and security becomes increasingly important (Krause et al., 2021). Protecting generation infrastructure from attacks is essential for maintaining the reliability and integrity of the entire SG system.
@@ 4.1.1. Energy Forecasting
A key aspect of generation in SGs is the integration of energy forecasting, particularity for renewable energy sources. SGs are designed to accommodate and efficiently manage the intermittent nature of renewable energy generation, such as solar and wind power (Wang et al., 2023). This integration helps reduce reliance on fossil fuels and promotes a more sustainable energy future. Specifically, energy forecasting leverages advanced machine learning techniques, including deep learning models like Long Short-Term Memory (LSTM) networks, to analyze historical data and predict future energy generation patterns (Malik and Pouttu, 2022).
Although accurate forecasting helps maintain grid stability by predicting fluctuations in renewable energy output, it also introduces new vulnerabilities, as attackers may target the prediction process. In the context of the FL framework, Shabbir et al. (Shabbir et al., 2024) focus on energy forecasting using FL in SG applications. The data involved typically consists of time-series data related to energy consumption and generation patterns. one primary attack method discussed is FDIA, which can take various forms, such as scaling attacks, random noise attacks, and mean-based attacks. These attacks aim to manipulate the learning process by injecting falsified data into the FL system. The impact of these attacks is often measured using metrics like Mean Absolute Percentage Error (MAPE). For instance, experiments have shown that scaling attacks can increase MAPE from 0.193% to 32.72%, while random noise attacks can elevate it from 0.183% to 129.75% as the proportion of attacked data rises from 10% to 100%. Badr et al. (Badr et al., 2022) propose a privacy-preserving FL approach for net-energy forecasting, addressesing the critical need for accurate energy predictions while protecting individual privacy in SG systems. To enhance privacy protection, researchers have introduced various strategies, including secure data aggregation (Fereidooni et al., 2021), differential privacy, and hybrid mechanisms. These approaches aim to mitigate potential privacy leaks that may occur through attacks on shared model parameters. FedGrid (Gupta et al., 2023) is an innovative framework that leverages FL to optimize energy distribution in SG systems. It addresses two major challenges in the energy sector: accurately predicting renewable energy generation and forecasting electric load, while simultaneously maintaining data privacy and security. By utilizing FL, FedGrid enables collaborative model training across multiple parties without compromising the privacy of individual data, a crucial factor when handling sensitive information like user electricity consumption patterns and localized weather data. FedGrid’s comprehensive analysis helps SG operators develop efficient supply strategies, optimizing the distribution of energy from both renewable and non-renewable sources.
@@
Generation Shabbir et al. (Shabbir et al., 2024) 2024 Energy Forecasting Time Series False Data Injection Security
Badr et al. (Badr et al., 2022) 2022 Net-Energy Forecasting Time Series Model Inversion, Membership Inference Privacy
FedGrid (Gupta et al., 2023) 2023 Net-Energy Forecasting Miscellaneous Data Inference Privacy
Moayyed et al. (Moayyed et al., 2022) 2022 Wind Speed Forecasting Miscellaneous Data Inference Privacy
T&D Lin et al. (Lin et al., 2022) 2022 State Estimation Measurements False Data Injection Security
Mohammadabadi et al. (Mohammadabadi et al., 2023) 2023 Event Classification Phasor Measurement Unit Data Inference Privacy
Jithish et al. (Jithish et al., 2023) 2023 Anomaly Detection Miscellaneous Data Integrity Security
Liu et al. (Liu and Wu, 2022) 2022 Voltage Control Miscellaneous Data Inference Privacy
Consumption Husnoo et al. (Husnoo et al., 2023) 2023 Load Forecasting Historical Data Integrity Security
Wen et al. (Wen et al., 2021) 2021 Energy Theft Detection Irish Smart Meter Inference, Model Poisoning Privacy, Security
Ashraf et al. (Ashraf et al., 2022) 2022 Energy Theft Detection Smart Meter Data Poisoning Security
Zafar et al. (Zafar et al., 2023) 2023 Energy Theft Detection Irish Smart Meter Data Poisoning Security
Li et al. (Li et al., 2024a) 2024 Energy Theft Detection Smart Meter Data Poisoning, Model Inversion Privacy, Security
@@ 4.1.2. Wind Speed Forecasting
Wind speed forecasting is a crucial process in renewable energy management, involving the prediction of future wind speeds using advanced statistical and machine learning techniques (Alhussein et al., 2019). This method utilizes historical wind speed data, along with other meteorological parameters, to generate accurate predictions for various time horizons, ranging from short-term (hours ahead) to long-term (days or weeks ahead) (Malik et al., 2022). Accurate wind speed forecasting directly contributes to improved wind power prediction, enabling the seamless integration of wind energy into the SG. By anticipating wind power output, grid operators can better optimize the balance between wind energy and other power sources, resulting in more efficient utilization of renewable resources (Ghofrani et al., 2014).
Wind speed forecasting often relies on advanced machine learning and deep learning algorithms (Lin et al., 2024; Sun et al., 2021), which can be vulnerable to adversarial attacks designed to manipulate their outputs. Attackers could exploit weaknesses in these algorithms, introducing errors in wind speed predictions and leading to inaccurate power generation forecasts. Within the FL framework, Moayyed et al. (Moayyed et al., 2022) propose a wind power forecasting framework integrated into the FL pipeline. The system predicts future wind power generation using multiple resource, including time-series data on wind power generation and meteorological data. Additionally, the model incorporates Sentinel-1 and Sentinel-2 satellite images for wind speed and bathymetry analysis, enhancing the spatial information available for forecasting. While the FL approach inherently provides some defense against data breaches by decentralizing sensitive information, local model updates from attackers can potentially compromise the entire system (Zhang et al., 2015; Liu et al., 2020).
@@ 4.2. Vulnerabilities in Transmission and Distribution Stages
SG technology revolutionizes traditional power grids by incorporating advanced ICTs into the existing energy infrastructure. This integration enables a more efficient, reliable, and flexible system for electricity transmission and distribution (Hamidi et al., 2010). The transmission component of the SG (Bayoumi, 2015) focuses on the high-voltage transport of electricity over long distances. In SGs, transmission systems benefit from advanced technologies such as Flexible AC Transmission Systems (FACTS) and High-Voltage Direct Current (HVDC) power transmission, which enhance the controllability, stability, and power transfer capability of AC transmission systems. However, transmission systems are vulnerable to various types of attacks. FDIAs are particularly concerning, as they can manipulate sensor measurements and mislead grid operators.
SG distribution networks (Konstantelos et al., 2016) are responsible for delivering electricity to end-users at lower voltages. These networks are undergoing significant changes to accommodate the increasing integration of DERs and renewable energy sources. Smart distribution systems enable bi-directional power flow, allowing consumers to also become producers of electricity. However, distribution systems face threats such as energy theft through meter tampering, privacy breaches of consumer data, and coordinated attacks on demand response systems (Ebtia et al., 2024).
@@ 4.2.1. State Estimation
AC State Estimation is a crucial process in SG systems, designed to determine the most likely steady-state operation of the power system based on field measurements. It serves as a fundamental data processing tool responsible for supporting and increasing system visibility while filtering errors that may appear in real-time measurements, system topology, and parameters (Acurio et al., 2023).
However, despite its significance, AC State Estimation is vulnerable to cyber threats, particularly FDIAs. These attacks can manipulate measurement data, introduce errors in state estimation, which can impact bus power flow calculations and the economic dispatch of the power system (Huang and Li, 2022). Within the FL framework, Lin et al. (Lin et al., 2022) propose an edge-based FL framework for detecting FDIAs in SG state estimation. This approach leverages local system operating data to predict and assess system stability in a distributed manner. By utilizing edge devices for local model training, the framework reduces the computational burden on central servers and enhances privacy protection. A key innovation in this work is the incorporation of an incentive mechanism designed to encourage high-quality data sharing and model training. This mechanism addresses the challenge of motivating participants to actively contribute to the FL process, which is essential for improving the overall detection accuracy of the system.
@@ 4.2.2. Events Classification and Anomaly Detection
Deep learning-based models (Niazazari and Livani, 2020), which have gained popularity for grid event classification, are particularly vulnerable to adversarial attacks. While highly effective under normal conditions, these models can be deceived by carefully crafted perturbations in the input data. The vulnerability arises from the complex nature of these models and their sensitivity to small changes in input features. Within the FL framework, Mohammadabadi et al. (Mohammadabadi et al., 2023) propose a federated XGBoost model to address data privacy concerns and enable collaborative learning. One key challenge in implementing FL for phasor measurement unit (PMU) data is the varying numbers of PMUs across different entities, resulting in diverse input structures for the event classification model. To overcome this, the proposed model constructs a common set of event features that can be generated by each entity, regardless of how many PMUs they possess. This study also considers two types of adversarial attacks on images using the Fast Gradient Sign Method (FGSM) and an iterative variant. In both approches, perturbations are applied to clean images to generate adversarial examples. For untargeted attacks, FGSM updates the adversarial image as follows:
@@
Jithish et al (Jithish et al., 2023) propose a FL-based solution that enables the training of machine learning models across multiple distributed devices without compromising user privacy or relying on extensive centralized computing resources. This approach not only preserves data confidentiality but also enables real-time anomaly detection, which is crucial for identifying and responding to various grid issues such as unusual power consumption patterns, infrastructure faults, and potential attacks. In this system, a global model is downloaded from the server to smart meters for on-device training. After local training, only the updated local model parameters are sent to the server to improve the global model, thus ensuring user privacy. To further protect the model parameter updates from potential adversaries, the SSL/TLS protocol is employed. adding an extra layer of security to guarantee the integrity and confidentiality of the transmitted data.

### v2

@@ 3.3. Consumption Stage
In the area of optimal power usage (rajesh2024federated; abdulla2024smart), FL facilitates the extraction of typical electricity consumption patterns from smart meter data, enabling more efficient grid management. For load forecasting (badr2023privacy; liu2023fedforecast; fekri2023asynchronous), FL-based methods have been developed to enhance the accuracy of short-term residential load predictions by utilizing distributed data from multiple sources. In the case of electricity theft detection (li2024fuse; alshehri2024deep; zafar2023step), frameworks like FedDetect employ federated learning to effectively identify fraudulent activities while preserving consumer privacy.
@@
Optimal Power Usage Rajesh et al. (rajesh2024federated) propose a FL framework designed to enhance privacy while optimizing power usage in SGs. The framework allows for collaborative learning across multiple participants without compromising individual user data. By securing power traces and enabling personalized recommendations for power usage optimization, the approach significantly improves privacy preservation in SG data analytics. This personalized aspect ensures that recommendations are tailored to the specific needs of users while maintaining data confidentiality. Similarly, Abdulla et al. (abdulla2024smart) present an adaptive FL approach for energy consumption forecasting using smart meter data. Their method improves forecasting accuracy by utilizing smart meter data without sharing raw information, ensuring user privacy. The adaptive nature of the approach allows it to adjust to individual consumption patterns, providing personalized forecasting models for each participant. This enhances both the accuracy of energy predictions and the privacy of users in the SG system.
@@
Load Forecasting Badr et al. (badr2023privacy) introduce a privacy-preserving and communication-efficient energy prediction scheme based on FL for SGs. Their approach enables accurate load forecasting while maintaining data confidentiality, which is crucial for infrastructure planning and power dispatching. By enhancing communication efficiency in the FL process, this method helps reduce power outages and equipment failures, making it an important tool for SG operations. Liu et al. (liu2023fedforecast) propose FedForecast, a FL framework designed for short-term probabilistic individual load forecasting. This framework provides comprehensive information about future load uncertainties while addressing data isolation and privacy concerns inherent in traditional centralized forecasting methods. By maintaining data confidentiality across households or smart meters, FedForecast improves the accuracy and reliability of short-term load forecasts. Fekri et al. (fekri2023asynchronous) develop an asynchronous adaptive FL algorithm for distributed load forecasting with smart meter data. This approach enhances the scalability and efficiency of load forecasting in SGs by adapting asynchronously to the data, thus addressing issues of privacy, security, and communication overhead found in traditional centralized methods. By doing so, the algorithm ensures accurate load forecasting while preserving the privacy of distributed data.
@@
Energy Theft Detection Li et al. (li2024fuse) introduce FUSE, a framework that combines FL with U-shape split learning for detecting electricity theft. FUSE identifies abnormal electricity consumption behaviors while preserving the privacy of both user data and the detection model. By integrating FL with split learning, FUSE enhances data protection and model security in smart grid systems, offering a robust solution for electricity theft detection. Alshehri et al. (alshehri2024deep) develop a deep anomaly detection framework that utilizes FL to detect electricity theft, including zero-day threats, in smart power grids. This framework preserves privacy while effectively identifying new and previously unseen attacks, thereby enhancing the security and stability of SGs. By addressing the challenge of detecting sophisticated electricity theft attempts, the framework plays a critical role in maintaining grid reliability. Zafar et al. (zafar2023step) present a FL-assisted hybrid deep learning model for electricity theft detection using smart meter data. This model enhances the security and reliability of SGs, particularly in the context of Industry 5.0. By leveraging FL to keep data decentralized, the model addresses privacy concerns while effectively detecting electricity theft. This hybrid approach offers a comprehensive solution to safeguard SGs from fraudulent activities.
@@ 4.1. Vulnerabilities in Generation Stage
In the context of SG, generation refers to the production of electricity from various sources, including both traditional and renewable energy (tuballa2016review). It is a crucial component of the SG pipeline, serving as the starting point for the flow of electricity throughout the entire system. As generation systems become more integrated and digitalized, privacy and security becomes increasingly important (krause2021cybersecurity). Protecting generation infrastructure from attacks is essential for maintaining the reliability and integrity of the entire SG system.
@@ 4.1.1. Energy Forecasting
A key aspect of generation in SGs is the integration of energy forecasting, particularity for renewable energy sources. SGs are designed to accommodate and efficiently manage the intermittent nature of renewable energy generation, such as solar and wind power (wang2023investigation). This integration helps reduce reliance on fossil fuels and promotes a more sustainable energy future. Specifically, energy forecasting leverages advanced machine learning techniques, including deep learning models like Long Short-Term Memory (LSTM) networks, to analyze historical data and predict future energy generation patterns (malik2022integration).
Although accurate forecasting helps maintain grid stability by predicting fluctuations in renewable energy output, it also introduces new vulnerabilities, as attackers may target the prediction process. In the context of the FL framework, Shabbir et al. (shabbir2024resilience) focus on energy forecasting using FL in SG applications. The data involved typically consists of time-series data related to energy consumption and generation patterns. one primary attack method discussed is FDIA, which can take various forms, such as scaling attacks, random noise attacks, and mean-based attacks. These attacks aim to manipulate the learning process by injecting falsified data into the FL system. The impact of these attacks is often measured using metrics like Mean Absolute Percentage Error (MAPE). For instance, experiments have shown that scaling attacks can increase MAPE from 0.193% to 32.72%, while random noise attacks can elevate it from 0.183% to 129.75% as the proportion of attacked data rises from 10% to 100%. Badr et al. (badr2022privacy) propose a privacy-preserving FL approach for net-energy forecasting, addressesing the critical need for accurate energy predictions while protecting individual privacy in SG systems. To enhance privacy protection, researchers have introduced various strategies, including secure data aggregation (fereidooni2021safelearn), differential privacy, and hybrid mechanisms. These approaches aim to mitigate potential privacy leaks that may occur through attacks on shared model parameters. FedGrid (gupta2023fedgrid) is an innovative framework that leverages FL to optimize energy distribution in SG systems. It addresses two major challenges in the energy sector: accurately predicting renewable energy generation and forecasting electric load, while simultaneously maintaining data privacy and security. By utilizing FL, FedGrid enables collaborative model training across multiple parties without compromising the privacy of individual data, a crucial factor when handling sensitive information like user electricity consumption patterns and localized weather data. FedGrid’s comprehensive analysis helps SG operators develop efficient supply strategies, optimizing the distribution of energy from both renewable and non-renewable sources.
@@
Generation Shabbir et al. (shabbir2024resilience) 2024 Energy Forecasting Time Series False Data Injection Security
Badr et al. (badr2022privacy) 2022 Net-Energy Forecasting Time Series Model Inversion, Membership Inference Privacy
FedGrid (gupta2023fedgrid) 2023 Net-Energy Forecasting Miscellaneous Data Inference Privacy
Moayyed et al. (moayyed2022cyber) 2022 Wind Speed Forecasting Miscellaneous Data Inference Privacy
T&D Lin et al. (lin2022incentive) 2022 State Estimation Measurements False Data Injection Security
Mohammadabadi et al. (mohammadabadi2023towards) 2023 Event Classification Phasor Measurement Unit Data Inference Privacy
Jithish et al. (jithish2023distributed) 2023 Anomaly Detection Miscellaneous Data Integrity Security
Liu et al. (liu2022federated) 2022 Voltage Control Miscellaneous Data Inference Privacy
Consumption Husnoo et al. (husnoo2023secure) 2023 Load Forecasting Historical Data Integrity Security
Wen et al. (wen2021feddetect) 2021 Energy Theft Detection Irish Smart Meter Inference, Model Poisoning Privacy, Security
Ashraf et al. (ashraf2022feddp) 2022 Energy Theft Detection Smart Meter Data Poisoning Security
Zafar et al. (zafar2023step) 2023 Energy Theft Detection Irish Smart Meter Data Poisoning Security
Li et al. (li2024fuse) 2024 Energy Theft Detection Smart Meter Data Poisoning, Model Inversion Privacy, Security
@@ 4.1.2. Wind Speed Forecasting
Wind speed forecasting is a crucial process in renewable energy management, involving the prediction of future wind speeds using advanced statistical and machine learning techniques (alhussein2019microgrid). This method utilizes historical wind speed data, along with other meteorological parameters, to generate accurate predictions for various time horizons, ranging from short-term (hours ahead) to long-term (days or weeks ahead) (malik2022multi). Accurate wind speed forecasting directly contributes to improved wind power prediction, enabling the seamless integration of wind energy into the SG. By anticipating wind power output, grid operators can better optimize the balance between wind energy and other power sources, resulting in more efficient utilization of renewable resources (ghofrani2014smart).
Wind speed forecasting often relies on advanced machine learning and deep learning algorithms (lin2024prediction; sun2021solar), which can be vulnerable to adversarial attacks designed to manipulate their outputs. Attackers could exploit weaknesses in these algorithms, introducing errors in wind speed predictions and leading to inaccurate power generation forecasts. Within the FL framework, Moayyed et al. (moayyed2022cyber) propose a wind power forecasting framework integrated into the FL pipeline. The system predicts future wind power generation using multiple resource, including time-series data on wind power generation and meteorological data. Additionally, the model incorporates Sentinel-1 and Sentinel-2 satellite images for wind speed and bathymetry analysis, enhancing the spatial information available for forecasting. While the FL approach inherently provides some defense against data breaches by decentralizing sensitive information, local model updates from attackers can potentially compromise the entire system (zhang2015byzantine; liu2020reflection).
@@ 4.2. Vulnerabilities in Transmission and Distribution Stages
SG technology revolutionizes traditional power grids by incorporating advanced ICTs into the existing energy infrastructure. This integration enables a more efficient, reliable, and flexible system for electricity transmission and distribution (hamidi2010smart). The transmission component of the SG (bayoumi2015power) focuses on the high-voltage transport of electricity over long distances. In SGs, transmission systems benefit from advanced technologies such as Flexible AC Transmission Systems (FACTS) and High-Voltage Direct Current (HVDC) power transmission, which enhance the controllability, stability, and power transfer capability of AC transmission systems. However, transmission systems are vulnerable to various types of attacks. FDIAs are particularly concerning, as they can manipulate sensor measurements and mislead grid operators.
SG distribution networks (konstantelos2016strategic) are responsible for delivering electricity to end-users at lower voltages. These networks are undergoing significant changes to accommodate the increasing integration of DERs and renewable energy sources. Smart distribution systems enable bi-directional power flow, allowing consumers to also become producers of electricity. However, distribution systems face threats such as energy theft through meter tampering, privacy breaches of consumer data, and coordinated attacks on demand response systems (ebtia2024spatial).
@@ 4.2.1. State Estimation
AC State Estimation is a crucial process in SG systems, designed to determine the most likely steady-state operation of the power system based on field measurements. It serves as a fundamental data processing tool responsible for supporting and increasing system visibility while filtering errors that may appear in real-time measurements, system topology, and parameters (acurio2023state).
However, despite its significance, AC State Estimation is vulnerable to cyber threats, particularly FDIAs. These attacks can manipulate measurement data, introduce errors in state estimation, which can impact bus power flow calculations and the economic dispatch of the power system (huang2022false). Within the FL framework, Lin et al. (lin2022incentive) propose an edge-based FL framework for detecting FDIAs in SG state estimation. This approach leverages local system operating data to predict and assess system stability in a distributed manner. By utilizing edge devices for local model training, the framework reduces the computational burden on central servers and enhances privacy protection. A key innovation in this work is the incorporation of an incentive mechanism designed to encourage high-quality data sharing and model training. This mechanism addresses the challenge of motivating participants to actively contribute to the FL process, which is essential for improving the overall detection accuracy of the system.
@@ 4.2.2. Events Classification and Anomaly Detection
Deep learning-based models (niazazari2020attack), which have gained popularity for grid event classification, are particularly vulnerable to adversarial attacks. While highly effective under normal conditions, these models can be deceived by carefully crafted perturbations in the input data. The vulnerability arises from the complex nature of these models and their sensitivity to small changes in input features. Within the FL framework, Mohammadabadi et al. (mohammadabadi2023towards) propose a federated XGBoost model to address data privacy concerns and enable collaborative learning. One key challenge in implementing FL for phasor measurement unit (PMU) data is the varying numbers of PMUs across different entities, resulting in diverse input structures for the event classification model. To overcome this, the proposed model constructs a common set of event features that can be generated by each entity, regardless of how many PMUs they possess. This study also considers two types of adversarial attacks on images using the Fast Gradient Sign Method (FGSM) and an iterative variant. In both approches, perturbations are applied to clean images to generate adversarial examples. For untargeted attacks, FGSM updates the adversarial image as follows:
@@
Jithish et al (jithish2023distributed) propose a FL-based solution that enables the training of machine learning models across multiple distributed devices without compromising user privacy or relying on extensive centralized computing resources. This approach not only preserves data confidentiality but also enables real-time anomaly detection, which is crucial for identifying and responding to various grid issues such as unusual power consumption patterns, infrastructure faults, and potential attacks. In this system, a global model is downloaded from the server to smart meters for on-device training. After local training, only the updated local model parameters are sent to the server to improve the global model, thus ensuring user privacy. To further protect the model parameter updates from potential adversaries, the SSL/TLS protocol is employed. adding an extra layer of security to guarantee the integrity and confidentiality of the transmitted data.